THE DETERMINATION OF C_{0} (OR A_{0}), D_{0}^{K}, H_{0}^{K}, AND SOME DARK STATES FOR SYMMETRIC-TOP MOLECULES FROM INFRARED SPECTRA WITHOUT THE NEED FOR LOACALIZED PERTURBA-TIONS

A. MAKI, 15012 24th Ave. S.E., Mill Creek, WA 98012-5718; \underline{T. MASIELLO}, National Institute of Standards and Technology, Gaithersburg, MD, 20899-8393; T.A. BLAKE , Pacific Northwest National Laboratory, P.O. Box 999, Mail Stop K8-88, Richland, WA 99352; J.W. NIBLER, Department of Chemistry, Oregon State University, Corvallis, OR 97331-4003; A. WEBER, National Institute of Standards and Technology, Gaithersburg, MD 20899-8393.

Several different combinations of allowed and apparently unperturbed rovibrational transitions are used to obtain the Kdependent rotational constants, C_{0}, D_{0}^{K}, and H_{0}^{K}. A necessary ingredient for the application of this technique is a band with selection rules Δk = ±1, Δl = ∓2, and appropriate hot or difference bands. Examples are given for boron trifluoride (BF_{3}), sulfur trioxide (SO_{3}), and cyclopropane (C_{3}H_{6}) for which there are microwave measurements that provide a check on the derived constants. Examples are also given for the determination of dark states from difference bands, and/or hot bands, and also whole forbidden bands that arise from mixing with distant energy levels.